6x^4 - 13x^2 + 5 = 0
Answer by mananth(16946)

6x^4 - 13x^2 + 5 = 0
6x^4-10x-3x^2+5=0
2x^2(3x^2-5)-1(3x^2-5)=0
(3x^2-5)(2x^2-1)=0
x= +/- sqrt(5/3)
OR
x=+/- sqrt(1/2)
...
m.ananth@hotmail.ca